Find the distance between (-1,7) and (-3,6).
a)10.07
b)13.15
c)15.28
d)2.23

Answer by checkley77(12844)

(-1,7) and (-3,6).
X distance=(-3+1)=2
Y distance=(7-6)=1
1^2+2^2=c^2
1+4=c^2
c^2=5
c=sqrt5
c=2.236 ans. D).